# Proper Body Landmark Subset Enables More Accurate and 5X Faster Recognition of Isolated Signs in LIBRAS

### v2

@@ 3.1.1 MINDS-Libras
This collection comprises 1,155 video recordings – each representing a sign – covering 20 signs selected according to phonological parameters (handshape, location, orientation, movement, and non-manual markers). The vocabulary includes “To happen”, “Student”, “Yellow”, “America”, “To enjoy”, “Candy”, “Bank”, “Bathroom”, “Noise”, “Five”, “To know”, “Mirror”, “Corner”, “Son”, “Apple”, “Fear”, “Bad”, “Frog”, “Vaccine”, and “Will”. The executions were performed by 12 signers varying in sex, age, and fluency in LIBRAS. Although clothing colors were not strictly standardized, most signers wore dark (black) attire, minimizing interference with the background.
Data collection was conducted in a controlled environment with a fixed chroma key background, and the sensors were positioned to capture the movement of the upper limbs. Two acquisition modalities were employed: an RGB camera (Canon EOS Rebel T5i, resolution 1920 $\times$ 1080) and an RGB-D sensor (Kinect v2, depth resolution 640 $\times$ 480), with only the RGB modality leveraged by our method. The typical recording distances were approx. 1.60m for the RGB-D sensor and approx. 1.92m for the RGB camera.
@@ 3.1.2 LIBRAS-UFOP
The recordings were performed by five signers (three female and two male). To increase variability, data collection took place in two different scenarios, under diverse illumination conditions, and without strict standardization of the distance between the participant and the sensor. In addition, subjects executed the signs at different speeds and considered both one-handed and two-handed executions, depending on the lexical nature of the sign. Data capturing was conducted at 30 fps, with resolution of 640 $\times$, 480 and a total of 3,040 sequences were obtained. A particular feature of this dataset is that each video comprises between 8 and 16 repetitions of the same sign performed by a given signer. Since the recognition pipeline requires each sign to be individualized, we use the same time cut points provided by Alves et al., (2024).
@@ 3.2 Recognition Performance Metrics
Recognition performance was measured in terms of accuracy, as well as macro-averaged precision, recall, and F1-score. These are traditional metrics in the ISLR domain (De Castro et al., , 2023; Alves et al., , 2024) derived from the counts of the confusion matrix in a multiclass context: TP (true positives), instances that belong to the target class and are correctly predicted as such; TN (true negatives), instances correctly predicted as not belonging to the target class; FP (false positives), instances that do not belong to the target class but are incorrectly predicted as belonging; and FN (false negatives), instances that belong to the target class but are incorrectly predicted as not belonging.
Accuracy quantifies the proportion of correct predictions over the total number of instances and serves as a global indicator of performance (Sokolova and Lapalme, , 2009; Powers, , 2011). Precision measures the reliability of positive predictions, while recall (sensitivity) measures the coverage of actual positives. To balance these aspects, the F1-score –the harmonic mean of precision and recall – often provides a more informative summary than accuracy in the presence of class imbalance. The use of precision, recall, and F1-score enables a more comprehensive evaluation and allows direct comparison with state-of-the-art methods, although no significant imbalance was observed in either evaluated dataset (see Figure 3).
Figure 3: Distribution of video sequence length w.r.t. signs for the evaluation datasets.
@@ 3.3 Experiments
The conducted experiments address one main research question and three additional ones:
@@
  • RQ3: How does our method, configured with the best-performing strategy, compare with existing approaches in the literature (Alves et al., , 2024; De Castro et al., , 2023; Cerna et al., , 2021; Passos et al., , 2021)?
  • RQ4: What speed-up does our MediaPipe-based solution provide compared to the OpenPose-based pipeline by Alves et al., (2024)?
@@ 3.3.1 Main Experiment (RQ1 + RQ2)
For each video, the full set of 543 landmarks was extracted using the MediaPipe Holistic pipeline. We retained the original raw data stream, containing the spatiotemporal trajectories of all points, as well as a processed stream obtained after the spline-based imputation (Section 2.3) to mitigate tracking errors and detection gaps. Note that both streams comprise the full set of detected landmarks, serving to establish a comparison baseline. From each stream, we derived four additional ones by applying the selection strategies introduced in Section 2.2 (Laines, Arcanjo, ASL-1st, and ASL-2nd), resulting, therefore, in 10 streams to be tested. These streams were encoded into $224\times 224\times 3$ skeleton images, following the method presented in Section 2.4.
Finally, each encoded image set was subjected to train-test sessions (Section 2.5). Following best practices (Alves et al., , 2024; De Castro et al., , 2023), we adopted a nested Leave-One-Person-Out (LOPO) cross-validation protocol, where data from one signer is used for testing while the remaining n1n-1 signers’ data are used for training and validation. In the nested approach, for a given signer in the test set, each of the n1n-1 remaining signers is detached once for validation, while the others are used for training. In total, n(n1)n(n-1) train–validation–test sessions were conducted.
@@ 3.3.2 Comparison with State-of-the-Art (RQ3)
This evaluation considered our skeleton-based approach with the ASL-2nd landmark selection strategy, which delivered the highest F1-score on both MINDS-Libras and LIBRAS-UFOP (last rows in Table 1). We then compared our results with Alves et al., (2024) – which inspired this work—, De Castro et al., (2023), Cerna et al., (2021), and the recent work of Naz et al., (2025). We also included a variation of the method by Alves et al., (2024), trained for 30 epochs instead of the original 20, ensuring fairer comparability with our proposal.
@@ 3.3.3 Speed-up Evaluation (RQ4)
This experiment compares the time performance of our approach – adopting the ASL-2nd landmark selection strategy – with Alves et al., (2024). While their method relies on OpenPose (Cao et al., , 2021) to extract the signer’s landmarks, our approach employs MediaPipe (Lugaresi et al., , 2019) for faster landmark extraction. For this evaluation, the video “02AlunoSinalizador08-5.mp4” from the MINDS-Libras dataset was selected, as its length (139 frames) is representative of the dataset’s average of 139.41 frames. Landmark extraction, the most time-consuming procedure, was measured for both approaches. Model inference was measured only for Alves et al., (2024), as the time difference was negligible.
@@ 3.4 Experimental Platform
The main experiment and comparison with state-of-the-art were run on a native Linux workstation (Ubuntu 22.04). The system features an Intel Core i9-10900KF at 3.70 GHz, 32 GB RAM, and an NVIDIA GeForce RTX 3060 GPU with 12 GB of VRAM. The software stack comprised Python 3.10.12 and PyTorch 2.0.1 with CUDA 12.1. The speed-up evaluation used an Intel Core i7-13700K at 5.4 GHz, with 32 GB of DDR4 RAM at 3200 MT/s, and an NVIDIA GeForce RTX 4070 GPU. Source code, landmarks data, and pretrained checkpoints will be released upon acceptance of this work.
Table 1: Results of the main experiment: comparative view of the landmark subsets. The last column reports the improvement (in percentage points) obtained with the spline-based imputation procedure. Values are presented as averages followed, in parentheses (when applicable), by the standard deviation.
Dataset Land. subset Accuracy Precision Recall F1-score F1-score imp. (p.p.)
MINDS-Libras All 0.70 (0.07) 0.71 (0.06) 0.69 (0.07) 0.66 (0.07) 6
Laines et al., (2023) 0.95 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.04) 4
Arcanjo et al., (2024) 0.95 (0.05) 0.96 (0.03) 0.94 (0.04) 0.94 (0.05) 4
ASL-1st 0.95 (0.05) 0.95 (0.05) 0.95 (0.05) 0.94 (0.05) 6
ASL-2nd 0.94 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.05) 4
LIBRAS-UFOP All 0.72 (0.04) 0.74 (0.02) 0.72 (0.04) 0.69 (0.04) 18
Laines et al., (2023) 0.88 (0.05) 0.90 (0.04) 0.88 (0.05) 0.87 (0.05) 17
Arcanjo et al., (2024) 0.91 (0.04) 0.92 (0.04) 0.91 (0.04) 0.90 (0.04) 5
ASL-1st 0.87 (0.05) 0.89 (0.04) 0.87 (0.05) 0.86 (0.05) 15
ASL-2nd 0.91 (0.04) 0.93 (0.03) 0.91 (0.04) 0.91 (0.04) 6
Values highlighted in bold indicate the highest average metric value for a given a dataset.
@@ 4 Results and Discussion
The following sections address, individually, the three experiments conducted in this work.
@@ 4.1 Main Experiment
Table 1 presents a comparative view of the recognition performance for the five landmark subsets considered in this work. For both datasets, using all landmarks provided by MediaPipe (All) yielded the worst results. For MINDS-Libras, the other landmark subsets performed very similarly, with small fluctuations in the average metrics and standard deviation values (within $\pm$2 p.p.). Nonetheless, a more pronounced contrast was observed for LIBRAS-UFOP. For instance, ASL-2nd achieved an F1-score of 0.91, which is 5 p.p. higher than ASL-1st and 4 p.p. higher than Laines et al., (2023). In summary, ASL-2nd and Arcanjo et al., (2024) presented the more consistent performance in terms of F1-score.
The reduction in input dimensionality (i.e., fewer landmarks) appears to enhance generalization while preserving the essential articulatory cues for isolated sign recognition. Unlike the baseline (All), the evaluated subsets focus primarily on pose and hand landmarks, with reduced or no reliance on dense facial meshes. This design helps mitigate noise caused by factors such as longer camera distances, illumination variability, and occasional occlusions—known to negatively affect facial landmark quality.
The last column in Table 1 shows the performance improvement (in terms of F1-score) resulting from applying the spline-based imputation procedure. Overall, the imputation enhanced the performance of all subsets by at least 4 p.p. The differences were more pronounced for LIBRAS-UFOP, where improvements of over 15 p.p. were observed for three subsets. The obtained results reinforce the need for post-processing the landmarks, as their detection with MediaPipe can be highly unstable.
@@
Table 2: Results of the comparative evaluation with state-of-the-art methods. Values are presented as averages followed, in parentheses (when applicable), by the standard deviation.
@@
Alves et al., (2024) (30 epochs) 2024 0.94 (0.04) 0.94 (0.05) 0.94 (0.04) 0.93 (0.04)
Alves et al., (2024) 2024 0.93 (0.05) 0.94 (0.05) 0.93 (0.05) 0.93 (0.05)
Naz et al., (2025) 2025 0.97 (0.01) - - -
De Castro et al., (2023) 2024 0.91 (0.07) - - 0.90 (0.08)
Passos et al., (2021) 2021 0.85 (0.02) - - -
@@
Alves et al., (2024) (30 epochs) 2024 0.82 (0.03) 0.84 (0.03) 0.82 (0.03) 0.80 (0.04)
Alves et al., (2024) 2024 0.82 (0.04) 0.83 (0.05) 0.81 (0.05) 0.80 (0.05)
Naz et al., (2025) 2025 0.89 (0.04) - - -
De Castro et al., (2023) 2024 0.74 (0.04) - - 0.71 (0.05)
Passos et al., (2021) 2021 0.65 (0.04) - - -
Cerna et al., (2021) 2021 0.74 (0.03) - - -
Values highlighted in bold indicate the highest average metric value for a given a dataset.
@@
4.2 Comparison with State-of-the-Art
Table 2 presents the results of the comparative evaluation with the literature. Results for Cerna et al., (2021) are reported only for LIBRAS-UFOP. On MINDS-Libras, our accuracy was comparable to Alves et al., (2024), who slightly benefited from training for 30 epochs instead of the original 20. Although Naz et al., (2025) achieved the highest accuracy, their evaluation relied on a less robust cross-validation protocol rather than LOPO.
On LIBRAS-UFOP, we outperformed the competing methods in all metrics, including Naz et al., (2025) (+2 p.p. in accuracy), which employed a less robust cross-validation protocol. Compared to Alves et al., (2024), our method achieved a substantially higher performance across all metrics (+11 p.p. in F1-score). These results demonstrate that a properly selected subset of landmarks, combined with an imputation procedure for missing landmarks, improves accuracy even when using a lighter framework for landmark extraction. The next section provides an alternative perspective on the gains, focusing on the speed-up achieved by our method.
4.3 Speed-up Evaluation
Table 3: Five runs of the recognition pipeline using a sample video from MINDS-Libras. The columns report the times for our approach and for Alves et al., (2024) (both in terms of landmark extraction), followed by the model inference time. The bottom row shows the mean (standard deviation).
# Ours Alves et al., (2024) Inf. time
@@ 4.3 Speed-up Evaluation
Table 3 presents the results of the speed-up evaluation. Notably, the average landmark extraction time in our pipeline is 4.4 seconds, compared to 28.7 seconds for the competing method, representing a speed-up nearly 6.7 times. When including the inference time, which better reflects the overall pipeline execution, the observed speed-up was approx. 5.6 times. This result, in conjunction with the accuracy improvement, reinforces the contribution of this work.
@@ 5 Conclusion and Future Work
In this paper, we investigated the feasibility of lightweight body landmark detection based on MediaPipe for the recognition of isolated signs in LIBRAS. Preliminary investigations using the pipeline proposed by Alves et al., (2024) indicated that simply replacing OpenPose with the MediaPipe Holistic model (using all available landmarks), although improving time efficiency, dramatically hindered recognition performance.
@@ ACKNOWLEDGEMENTS
Omitted due to anonymous submission.

### v3

@@ 3.1.1 MINDS-Libras
It comprises 1,155 video recordings – each representing a sign – covering 20 signs selected according to phonological parameters (handshape, location, orientation, movement, and non-manual markers). The executions were performed by 12 signers varying in sex, age, and fluency in LIBRAS. Although clothing colors were not strictly standardized, most signers wore dark (black) attire, minimizing background cluttering. Data was collected in a controlled environment with a fixed chroma key background and sensors positioned to capture the movement of the upper limbs.
Two acquisition modalities were employed: an RGB camera (Canon EOS Rebel T5i, resolution 1920 $\times$ 1080) and an RGB-D sensor (Kinect v2, depth resolution 640 $\times$ 480), with only the RGB modality leveraged in this study. The typical recording distances were (approximately) 1.60m for the RGB-D sensor and 1.92m for the RGB camera.
@@ 3.1.2 LIBRAS-UFOP
The recordings were performed by five signers in two different scenarios, under diverse illumination conditions, and without strict standardization of the distance between the participant and the sensor. The signs were executed at different speeds and considered both one- and two-handed executions, depending on the lexical nature of the sign. Recordings were captured at 30 fps, with a resolution of 640 $\times$ 480, totalling 3,040 video sequences. The original recordings comprise between 8 and 16 repetitions (in sequence) of the same sign for a given signer. For consistency with the ISLR framework, which requires a single sign execution per video, we use the same time cut points as [1].
@@ 3.2 Recognition Performance Metrics
Performance was measured in terms of traditional metrics in the ISLR domain [6, 1]: accuracy, as well as macro-averaged precision, recall, and F1-score. Accuracy quantifies the proportion of correct predictions over the total number of instances and serves as a global indicator of performance. Precision measures the reliability of positive predictions, while recall (sensitivity) measures the coverage of actual positives. To balance these aspects, the F1-score – the harmonic mean of precision and recall – often provides a more informative summary than accuracy in the presence of class imbalance.
@@ 3.3 Experiments
The conducted experiments address four research question:
@@
  • RQ3: How does our method, configured with the best-performing strategy, compare with existing approaches in the literature [1, 6, 5, 16]?
  • RQ4: What speed-up does our MediaPipe-based solution provide compared to the OpenPose-based pipeline by Alves et al. [1]?
@@ 3.3.1 Main Experiment (RQ1 + RQ2)
For each video, the full set of 543 landmarks was extracted with the MediaPipe Holistic pipeline. We retained the original raw data stream, containing the spatiotemporal trajectories of all points, as well as a processed stream obtained after imputation (Section 2.3) to mitigate detection or tracking failures. Note that both streams comprise the full set of detected landmarks, serving to establish a comparison baseline. From each stream, four additional were derived by applying the selection strategies introduced in Section 2.2 (Laines, Arcanjo, ASL-1st, and ASL-2nd), resulting in 10 streams, later encoded into $224\times 224\times 3$ skeleton images (Section 2.4).
Finally, each encoded image set was subjected to train-test sessions (Section 2.5). Following best practices [1, 6], we adopted a nested Leave-One-Person-Out (LOPO) protocol, where data from one signer is used for testing while the remaining n1n-1 signers’ data is used for training and validation. In the nested approach, each of the n1n-1 signers is detached once for validation, while the others are used for training. In total, n(n1)n(n-1) train–validation–test sessions were conducted.
@@ 3.3.2 Comparison with State-of-the-Art (RQ3)
This evaluation focused on the ASL-2nd landmark selection strategy, which delivered the highest F1-score on both MINDS-Libras and LIBRAS-UFOP (last rows in Table 1). Our results were compared with Alves et al. [1] – which inspired this work–, De Castro et al. [6], Cerna et al. [5], and the recent work of Naz et al. [15]. We also tested [1] with 30 epochs (rather than 20) for fairer comparability.
@@ 3.3.3 Speed-up Evaluation (RQ4)
This experiment compares the time performance of our approach – adopting the ASL-2nd landmark selection strategy – with [1]. While they rely on OpenPose [4] to extract landmarks, our approach employs MediaPipe [12] for faster landmark extraction. In this evaluation, the video “02AlunoSinalizador08-5.mp4” from the MINDS-Libras dataset was selected, as its length (139 frames) is representative of the dataset’s average of 139.41 frames. Landmark extraction, the most time-consuming procedure, was measured for both approaches. Model inference was measured only for [1] since the same CNN model is employed in both cases.
@@ 3.4 Experimental Platform
Speed-up evaluation: Intel Core i7-13700K at 5.4 GHz, with 32 GB of DDR4 RAM at 3200 MT/s, and a NVIDIA GeForce RTX 4070 GPU. Other experiments: Intel Core i9-10900KF at 3.70 GHz, 32 GB RAM, and a NVIDIA GeForce RTX 3060 GPU. The software [^1] Source code, landmarks data, and pretrained checkpoints will be released upon acceptance of this work. stack comprised Python 3.10 and PyTorch 2.0.1 with CUDA 12.1.
@@ 4.1 Main Experiment
Table 1 shows a comparative view for the five landmark subsets considered in this work. For both datasets, the baseline (All) yielded the worst results. For MINDS-Libras, the four landmark subsets performed very similarly, with small fluctuations in the average and standard deviation values (within $\pm$ 2 p.p.). A more pronounced contrast was observed for LIBRAS-UFOP: ASL-2nd achieved an F1-score of 0.91, which is 5 p.p. higher than ASL-1st and 4 p.p. higher than Laines [11]. In summary, ASL-2nd and Arcanjo [2] presented a more consistent F1-score.
Table 1: Comparing landmark subsets (main experiment).
Land. subset Accuracy Precision Recall F1-score F1 imp.
MINDS-Libras
All 0.70 (0.07) 0.71 (0.06) 0.69 (0.07) 0.66 (0.07) 6
Laines[11] 0.95 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.04) 4
Arcanjo [2] 0.95 (0.05) 0.96 (0.03) 0.94 (0.04) 0.94 (0.05) 4
ASL-1st [20] 0.95 (0.05) 0.95 (0.05) 0.95 (0.05) 0.94 (0.05) 6
ASL-2nd [23] 0.94 (0.04) 0.95 (0.05) 0.95 (0.04) 0.94 (0.05) 4
LIBRAS-UFOP
All 0.72 (0.04) 0.74 (0.02) 0.72 (0.04) 0.69 (0.04) 18
Laines [11] 0.88 (0.05) 0.90 (0.04) 0.88 (0.05) 0.87 (0.05) 17
Arcanjo [2] 0.91 (0.04) 0.92 (0.04) 0.91 (0.04) 0.90 (0.04) 5
ASL-1st [20] 0.87 (0.05) 0.89 (0.04) 0.87 (0.05) 0.86 (0.05) 15
ASL-2nd [23] 0.91 (0.04) 0.93 (0.03) 0.91 (0.04) 0.91 (0.04) 6
@@
Average values followed by the standard deviation in parentheses (when applicable).
The last column reports the improvement (p.p.) yielded by spline-based imputation.
Values in bold indicate the highest average metric value for a given dataset.
The reduction in input dimensionality (i.e., fewer landmarks) benefits generalization while preserving the essential articulatory cues for isolated sign recognition. Unlike the baseline (All), the other subsets focus primarily on pose and hand landmarks, with reduced or no reliance on dense facial meshes. This design helps mitigate noise caused by factors such as longer camera distances, illumination variability, and occasional occlusions that affect facial landmark quality.
The last column in Table 1 shows the F1-score improvement (in percent points) yielded by spline-based imputation (e.g., F1-score for Laines increases by 4 p.p. through imputation). Overall, imputation enhanced the performance of all subsets by at least 4 p.p. The differences were more pronounced for LIBRAS-UFOP, where improvements of over 15 p.p. were observed in three cases. The obtained results reinforce the need for post-processing the landmarks, as their detection with MediaPipe can be highly unstable.
4.2 Comparison with State-of-the-Art
Table 2 shows the results of the comparative evaluation with the literature. Results for [5] are reported only for LIBRAS-UFOP. On MINDS-Libras, our accuracy was comparable to [1], which slightly benefited from increasing training (30 epochs rather than 20). Although [15] achieved the highest accuracy, their evaluation relied on a less robust evaluation protocol than LOPO. For reference, De Castro et al. [6] observed an accuracy drop of nearly 40 p.p. when switching from traditional $k$-fold to LOPO, highlighting how strongly results depend on the evaluation protocol.
On LIBRAS-UFOP, we outperformed the competing methods in all metrics, including [15] (+2 p.p. in accuracy), which did not employ LOPO. Compared to [1], we achieved significative higher performance across all metrics (+11 p.p. in F1-score). These results reveal that a proper subset of landmarks, combined with an imputation procedure, is highly beneficial even when using lighter landmark extraction.
4.3 Speed-up Evaluation
Table 3 presents the results of the speed-up evaluation. Notably, the average landmark extraction time in our pipeline is 4.4 seconds, compared to 28.7 seconds for the competing method, representing a speed-up of nearly 6.7 times. When including the inference time, which better reflects the overall pipeline execution, the observed speed-up was approximately 5.6 times. It is worth noting that strict real-time performance depends on the deployment configuration, such as the hardware setup. Although this was not our focus, the obtained results reveal an advance toward real-time ISLR.
Table 2: Comparison with the State-of-the-Art.
@@ 4.3 Speed-up Evaluation
Alves et al. [1] (30 epochs) 2024 0.94 (0.04) 0.94 (0.05) 0.94 (0.04) 0.93 (0.04)
Alves et al. [1] 2024 0.93 (0.05) 0.94 (0.05) 0.93 (0.05) 0.93 (0.05)
Naz et al. [15] 2025 0.97 (0.01) - - -
De Castro et al. [6] 2024 0.91 (0.07) - - 0.90 (0.08)
Passos et al. [16] 2021 0.85 (0.02) - - -
@@
Alves et al. [1] (30 epochs) 2024 0.82 (0.03) 0.84 (0.03) 0.82 (0.03) 0.80 (0.04)
Alves et al. [1] 2024 0.82 (0.04) 0.83 (0.05) 0.81 (0.05) 0.80 (0.05)
Naz et al. [15] 2025 0.89 (0.04) - - -
De Castro et al. [6] 2024 0.74 (0.04) - - 0.71 (0.05)
Passos et al. [16] 2021 0.65 (0.04) - - -
Cerna et al. [5] 2021 0.74 (0.03) - - -
Average values followed by the standard deviation in parentheses (when applicable). Values in bold indicate the highest average metric value for a given dataset
@@
Table 3: Five runs of the ISLR framework using a test case from MINDS-Libras.
# Land. time (Ours) Land. time [1] Inf. time
@@
The bottom row shows the mean (standard deviation).
@@ 5 Conclusion and Future Work
In this paper, we investigated the feasibility of lightweight body landmark detection with MediaPipe for the recognition of isolated signs in LIBRAS. Preliminary investigation using the ISLR framework in [1] indicated that simply replacing OpenPose with the MediaPipe Holistic model (using all available landmarks), although improving time efficiency, dramatically hindered recognition performance.
@@ Acknowledgements
We acknowledge the support of FAPES and CAPES (Process 2021-2S6CD, FAPES No. 132/2021) through the PDPG (Programa de Desenvolvimento da Pós-Graduação – Parcerias Estratégicas nos Estados).